# Quality-Driven Curation of Remote Sensing Vision-Language Data via Learned Scoring Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have demonstrated great potential in interpreting remote sensing (RS) images through language-guided semantic understanding. However, the effectiveness of these VLMs critically depends on high-quality image-text training data that captures rich semantic relationships between visual content and language descriptions. Unlike natural images, RS lacks large-scale interleaved image-text pairs from web data, making data collection challenging. While current approaches rely primarily on rule-based methods or flagship VLMs for data synthesis, a systematic framework for automated quality assessment of such synthetically generated RS vision-language data is notably absent. To fill this gap, we propose a novel score model trained on large-scale RS vision-language preference data for automated quality assessment. Our empirical results demonstrate that fine-tuning CLIP or advanced VLMs (e.g., Qwen2-VL) with the top 30% of data ranked by our score model achieves superior interpretation accuracy compared to both full-data fine-tuning and CLIP-score-based ranking approaches. Furthermore, we demonstrate applications of our scoring model for reinforcement learning (RL) training and best-of-N (BoN) test-time scaling, enabling significant improvements in VLM performance for RS tasks.
@@ 1 Introduction
Figure 1: Examples from common RS vision-language datasets highlighting quality issues. These samples illustrate three prevalent problems: uninformative descriptions that lack meaningful content about the image features, hallucinated answers that introduce factual inaccuracies, and responses to infeasible questions that cannot be answered based on the visible images
@@
The advancement of artificial intelligence has consistently demonstrated the benefits of scaling across three crucial dimensions: data volume, computational resources, and model complexity [16, 12]. In visual understanding, vision-language models (VLMs) have particularly benefited from diverse training data, exhibiting evolutionary patterns that mirror human cognitive development in environmental perception and interpretation. From the foundational CLIP model [29] to recent integrations with large language models (LLMs) [5, 38, 20, 8], VLMs have achieved remarkable success, emerging as indispensable tools for real-world applications ranging from intelligent agents [52] to autonomous driving systems [26].
@@
Despite these advancements, VLMs demonstrate limited capability in interpreting remote sensing (RS) images due to substantial domain shifts in data distributions [19, 17]. This performance gap stems primarily from a critical bottleneck: the scarcity of large-scale vision-language data for RS domains, which, unlike natural images, lacks the abundant, naturally aligned image-text pairs readily available through web crawling. Current approaches to bridge this gap include: (1) binary classification models filtering RS data from general vision-language datasets like DataComp [51], (2) rule-based construction of vision-language pairs using OpenStreetMap tags [39], and (3) automated data generation through flagship VLMs [22, 17, 24, 18, 27]. While these methods advance holistic RS understanding, they introduce significant quality concerns: rule-based approaches often produce informationally sparse or semantically inconsistent pairs, while VLM-generated data risks propagating hallucinatory content that misrepresents the actual imagery and incorrectly infers answers to questions that cannot be determined from the given images (Fig. 1). The absence of robust, automated quality assessment framework consequently constrains further improvements in RS-specific VLMs [18]
@@
To address this critical gap, we propose ScoreRS – a learned quality scoring model for vision-language data in RS domains. In developing ScoreRS, we first construct both image-caption and vision instruction preference datasets by leveraging existing RS-specific VLMs and robust open-source VLMs as policy models, while employing both rule-based evaluation and flagship VLMs as judges for preference dataset construction. Through a three-stage progressive training process with the constructed data, we evaluate ScoreRS’s effectiveness in both data quality ranking and its practical applications as a large VLM reward model for GRPO [33] training and best-of-n (BoN) selector. Our results demonstrate that fine-tuning CLIP and Qwen2VL [38] on just 30% of the data, ranked using ScoreRS scores, outperforms models trained on either the complete dataset or CLIP score-ranked data. Furthermore, evaluation on the challenging RS-specific vision-language benchmark VG-DOIR [47] and LHRS-Bench [24] reveals that ScoreRS serves as an effective BoN selector for improving results when scaling VLMs with multiple generated samples at test time, and can be integrated with rule-based rewards for GRPO reinforcement training to enhance model capabilities.
@@
    We introduce the first large-scale RS-specific preference dataset, comprising pairwise preference pairs for both image-captions and vision instructions, alongside ScoreRS—a novel data scoring model specifically designed for quality control of RS vision-language data.
@@
    We demonstrate that models fine-tuned on RS vision-language data ranked by ScoreRS scores consistently outperform those trained on either the complete dataset or CLIP score-ranked data.
@@
    We empirically validate ScoreRS’s effectiveness through two key applications: (1) as a reward model for GRPO training, and (2) as a BoN selector for scaling VLM-generated samples at test time. Both applications yield significant improvements in RS VLM performance across multiple challenging benchmarks.
@@ 2 Method
Figure 2: Our data pipeline for generating pairwise preference datasets and the training and downstream application of our ScoreRS model. $I_{i}\in\mathcal{I}$ represents a RS image, and $T_{i}\in\mathcal{T}$ represents an image caption, question, or conversation associated with the image
@@
Formally, we define a scoring function $f_{\theta}:\mathcal{T}\times\mathcal{I}\rightarrow\mathbb{R}$, parameterized by $\theta$, where $\mathcal{T}$ and $\mathcal{I}$ denote the text and image spaces, respectively. Given an image-text pair $(I,T)\in\mathcal{I}\times\mathcal{T}$, which may consist of either a simple caption or a multi-turn conversation associated with the image, the model outputs a scalar quality score $s=f_{\theta}(I,T)$. The scoring function is trained to assign higher scores to better-aligned image-text pairs $(I,T)$ through pairwise preference learning. For each image $I$, we collect pairs of text $(T^{+},T^{-})$ where $T^{+}$ is preferred over $T^{-}$. This preference dataset is formally defined as $\mathcal{D}={(I_{i},T^{+}_{i},T^{-}_{i})}_{i=1}^{|\mathcal{D}|}$. The pairwise preferences are modeled using the Bradley-Terry model [2], which defines the probability of $T^{+}$ being preferred over $T^{-}$ given image $I$ as:
$p(T^{+}\succ T^{-}|I)=\frac{\exp(f_{\theta}(I,T^{+}))}{\exp(f_{\theta}(I,T^{+}))+\exp(f_{\theta}(I,T^{-}))},$ (1)
where $\succ$ denotes the preference relation. The model parameters $\theta$ are optimized by minimizing the empirical negative log-likelihood loss [25, 37]:
@@
Given the absence of RS specific vision-language preference datasets, we start by establishing a data collection pipeline for acquiring both image-caption and vision instruction preference pairs. Subsequently, we detail the construction of our scoring function $f_{\theta}$ by leveraging pre-trained VLMs and employing a progressive training strategy.
@@ 2.1.1 Image-Caption Preference Dataset
Following established practices in VLM training [38, 24, 27, 20], we start with training ScoreRS to judge high-quality image-caption alignments. Since RS images exhibit variations across geographical locations [24, 39], we leverage the LHRS-Align dataset [24] as our image source. This dataset comprises 1.15M orthorectified RS images from major global urban areas, enabling ScoreRS to learn from diverse geographical contexts worldwide.
Deduplication
We implement a rigorous image deduplication process to further ensure image quality and representativeness: (1) Feature extraction using the SSCD copy detection model [28, 10] to compute image embeddings. (2) Similarity computation through cosine distance computation between image embeddings, with a carefully tuned threshold of 0.65 (empirically determined through experiments across the range 0.6-0.9). (3) Duplicate grouping via connected-components algorithm, preserving one image-text pair per component. Finally, this deduplication pipeline yields 76K distinctive representative RS images.
Pairwise Preference Generation
Our caption pairwise preference generation process employs three VLMs: the RS-specific LHRS-Bot-Nova [18] and two general-purpose models, Qwen2VL-7B [38] and InternVL-2.5-8B [5]. We generate captions for each image using same prompts and sampling parameters across all models. The resulting captions are then evaluated by GPT-4o along five dimensions: accuracy, completeness, conciseness, objectivity, and spatial clarity. For each image, we compute the mean score across these five dimensions for each caption. The caption with the highest average score is designated as the positive example ($T^{+}$), while the remaining two captions serve as negative examples ($T^{-}$). After removing results with parsing errors, we construct a dataset containing 72K image-caption preference pairs. To validate the reliability of GPT-4o as a preference judge, we conducted a human evaluation study on a random sample of 1,000 pairs. Human experts independently annotated these samples to identify positive and negative captions. The results showed 92.6% agreement (926 out of 1,000 samples) between human judgments and GPT-4o’s assessments. This high level of agreement validates our choice of GPT-4o as a reliable judge for generating pairwise preferences.
@@ 2.1.2 Vision Instruction Preference Dataset
Vision instruction data consists of conversations about images [20]. To construct a RS-specific vision instruction preference dataset that enables ScoreRS to evaluate responses to diverse user queries, we prioritized collecting a broad spectrum of question types and conversation scenarios. We aggregate vision instruction data from multiple sources: GeoChat [17] (306K samples), LHRS-Instruct [24] (39.8K samples), and a subset of SkysenseGPT [22] (381K samples) as the source of our vision instruction preference dataset.
Deduplication
We employ a two-stage similarity-based filtering process to ensure question diversity: (1) Text similarity: let $Q=\{T_{1},T_{2},\ldots,T_{n}\}$ be the set of all questions extracted from the conversations. We compute their semantic embeddings using the BGE [3] model, where for any two queries $T_{i},T_{j}\in Q:\text{sim}_{\text{text}}(T_{i},T_{j})=\cos(\text{BGE}(T_{i}),\text{BGE}(T_{j}))$; (2) Image similarity: For query pairs where $\text{sim}_{\text{text}}(T_{i},T_{j})>0.65$, we further examine their corresponding images $I_{i},I_{j}$ using SSCD [28] embeddings: $\text{sim}_{\text{image}}(I_{i},I_{j})=\cos(\text{SSCD}(I_{i}),\text{SSCD}(I_{j}))$. When both $\text{sim}_{\text{text}}(T_{i},T_{j})$ and $\text{sim}_{\text{image}}(I_{i},I_{j})$ exceed 0.65, we retain only one question-image sample from the pair to minimize redundancy.
Pairwise Preference Generation
@@
After obtaining the deduplicated set of conversations, we categorized them into two types: close-ended questions with definitive answers (e.g., vision-question answering, classification, and visual grounding tasks) and open-ended questions requiring interpretive responses (e.g., visual reasoning). For close-ended questions, which can be objectively verified against specific correct responses, we prompted the Qwen2VL-7B model to generate an answer $\hat{y}$. If $\hat{y}$ differs from the provided answer $y$, we extract the sample into our vision instruction preference dataset, treating the generated answer as the negative target $T^{-}$ and the provided answer as the positive target $T^{+}$. For open-ended questions, we first employ LHRS-Bot-Nova and Qwen2VL-7B to generate answers $\hat{y}_{1}$ and $\hat{y}_{2}$, respectively, given the input image and question. Subsequently, we prompt Qwen2VL-72B to evaluate the generated answers $\hat{y}_{1}$ and $\hat{y}_{2}$, as well as the answer $y$ from the source dataset, across five dimensions: accuracy, completeness, clarity, relevance, and hallucination. Qwen2VL-72B assigns scores (0-5) for each dimension and provides an overall confidence score (0-5) for the evaluation. We select samples where the confidence score exceeds 3 and designate the answer with the highest average score across the five evaluation dimensions as the positive target ($T^{+}$). Both remaining answers are treated as negative targets ($T^{-}$), which simultaneously enhances the diversity of negative samples and increases our training data volume. We do not directly consider the answer from the source dataset as $T^{+}$ because many of them are generated by older VLMs and are often too short, meaningless, or contain incorrect statements. To manage costs associated with the large volume of data, we employ Qwen2VL-72B rather than GPT-4o as our evaluator. After filtering out parsing errors and applying human-defined quality rules, we curate a dataset of 26K vision instruction preference pairs. The detailed prompts used in this process are provided in Sec. 7.2.
@@
During the process of collecting the question set $Q$, we observe that most data sources lack RS application-specific questions (e.g., questions related to agricultural and disaster analysis). To enhance ScoreRS’s performance on domain-specific content, we systematically develop a specialized question set comprising five manually crafted questions for each of 12 expert-defined RS image analysis categories. We then randomly sample 35K RS images from our deduplicated image dataset. For each image, we randomly select a question from the manually designed question set and prompt Qwen2.5-13B to rephrase the question, increasing the diversity of the manually generated questions. Subsequently, we employ LHRS-Bot-Nova and Qwen2VL-7B to generate answers given the image and rephrased question. Finally, we use the same prompt to elicit GPT-4o to evaluate the generated answers and select the highest-scoring answer as the positive target $T^{+}$, while treating the remaining answers as negative targets $T^{-}$. This process yield 33K RS-specific vision instruction preference data points after removing entries with parsing errors. The manually designed questions and categories can be found in Sec. 7.2.

### v2

@@ Abstract
Vision-Language Models (VLMs) have demonstrated great potential in interpreting remote sensing (RS) images through language-guided semantic. However, the effectiveness of these VLMs critically depends on high-quality image-text training data that captures rich semantic relationships between visual content and language descriptions. Unlike natural images, RS lacks large-scale interleaved image-text pairs from web data, making data collection challenging. While current approaches rely primarily on rule-based methods or flagship VLMs for data synthesis, a systematic framework for automated quality assessment of such synthetically generated RS vision-language data is notably absent. To fill this gap, we propose a novel score model trained on large-scale RS vision-language preference data for automated quality assessment. Our empirical results demonstrate that fine-tuning CLIP or advanced VLMs (e.g., Qwen2-VL) with the top 30% of data ranked by our score model achieves superior accuracy compared to both full-data fine-tuning and CLIP-score-based ranking approaches. Furthermore, we demonstrate applications of our scoring model for reinforcement learning (RL) training and best-of-N (BoN) test-time scaling, enabling significant improvements in VLM performance for RS tasks. Our code, model, and dataset are publicly available111 [Uncaptioned image] Data [Uncaptioned image] Model [Uncaptioned image] Code .
@@ 1 Introduction
Figure 1: Examples from RS vision-language datasets showing quality issues across five dimensions. Green represents reasonably good expression, while red represents low-quality expression
@@
The advancement of artificial intelligence has consistently benefited from scaling across three crucial dimensions: data volume, computational resources, and model complexity (kaplan2020scaling, ; hoffmann2022training, ). In visual understanding, vision-language models (VLMs) have particularly benefited from diverse training data, exhibiting evolutionary patterns that mirror human cognitive development in perception and interpretation. From the foundational CLIP model (radford2021learning, ) to recent integrations with large language models (LLMs) (chen2024expanding, ; wang2024qwen2, ; liu2024visual, ; dong2025scalable, ), VLMs have achieved remarkable success, emerging as indispensable tools for real-world applications ranging from agents system (zheng2024gpt, ) to autonomous driving (pan2024vlp, ).
@@
Despite these advancements, VLMs demonstrate limited capability in interpreting remote sensing (RS) images due to substantial domain shifts in data distributions (liu2024remoteclip, ; kuckreja2024geochat, ). This performance gap stems primarily from a critical bottleneck: the scarcity of large-scale vision-language data for RS domains, which, unlike natural images, lacks the abundant, naturally aligned image-text pairs readily available through web crawling. Current approaches to bridge this gap include: (1) binary classification models filtering RS data from general vision-language datasets like DataComp (zhang2023rs5m, ), (2) rule-based construction of vision-language pairs using OpenStreetMap tags (wang2024skyscript, ), and (3) automated data generation through flagship VLMs (luo2024skysensegpt, ; kuckreja2024geochat, ; muhtar2024lhrs, ; li2024lhrs, ; pang2024vhmversatilehonestvision, ) or leveraging supervision from internet images (mall2023remote, ). While these methods advance holistic RS understanding, they introduce significant quality concerns: rule-based approaches often produce informationally sparse or semantically inconsistent pairs, while VLM-generated data risks propagating hallucinatory content that misrepresents the actual imagery and incorrectly infers answers to questions that cannot be determined from the given images (Figure 1). The absence of robust, automated quality assessment framework consequently constrains further improvements in RS-specific VLMs (li2024lhrs, )
@@
To address this critical gap, our method commences with the definition of text quality for RS images, outlining five crucial dimensions that characterize high-quality descriptions or instruction samples: relevance to visual content (pang2024vhmversatilehonestvision, ; luo2024skysensegpt, ; xue2024enhanced, ), specificity and detail (li2024lhrs, ; li2024if, ; zhou2025benchmark, ), completeness of coverage for salient features (chan2023clair, ; zhou2025benchmark, ), clarity and fluency of language (wang2024skyscript, ), and semantic richness for RS applications (wen2025rs, ; silva2402large, ). Guided by these quality dimensions, we construct preference datasets for both image-captions and vision instructions by employing a diverse array of policy models—including RS-specialized VLMs and open-source alternatives—while utilizing a combination of rule-based evaluation metrics and flagship VLMs as objective judges. Building upon these curated preference datasets, we introduce ScoreRS, a novel learned quality scoring model tailored for RS vision-language data, developed via a three-stage progressive training recipe. We evaluate ScoreRS’s effectiveness in both data quality ranking and its practical applications as a large VLM reward model for group relative policy optimization (GRPO) (shao2024deepseekmath, ) reinforcement learning (RL) and best-of-n (BoN) selector. Our results demonstrate that fine-tuning CLIP and Qwen2VL (wang2024qwen2, ) on just 30% of the data, ranked using ScoreRS scores, outperforms models trained on either the complete dataset or CLIP-score ranked data. Furthermore, evaluation on the challenging RS-specific vision-language benchmark VG-DIOR (zhan2023rsvg, ) and LHRS-Bench (muhtar2024lhrs, ) reveals that ScoreRS can be integrated with rule-based rewards for GRPO RL to enhance model capabilities, and serves as an effective BoN selector for improving results when scaling VLMs with multiple generated samples at test time.
@@
    We establish a framework for defining text quality in RS vision-language data and introduce the first large-scale RS-specific preference dataset, comprising pairwise preference pairs for both image-captions and vision instructions. Based on this dataset, we develop ScoreRS—a novel data scoring model specifically designed for automated quality assessment of RS vision-language data.
@@
    We demonstrate that models fine-tuned on just the top 30% of RS vision-language data ranked by ScoreRS scores consistently outperform those trained on either the complete dataset or data ranked using CLIP scores, highlighting the efficacy of our quality-driven curation approach.
@@
    We validate ScoreRS’s effectiveness through two practical applications: (1) as a reward model for GRPO-based RL, and (2) as a BoN selector for VLM-generated samples at test time. Both applications yield improvements in RS VLM performance across multiple challenging benchmarks.
2 Related Work
2.1 RS Vision-Language Data Curation
Unlike the general vision domain, where vision-language data can be readily crawled from abundant image-text interleaved webpages (xu2023demystifying, ; gadre2023datacomp, ; tong2024cambrian, ; schuhmann2022laion, ; chen2024sharegpt4v, ), the RS domain presents a challenge. While rich in open-source image data (muhtar2024lhrs, ; muhtar2023cmid, ; tamiminia2020google, ), RS lacks corresponding text descriptions and analytical conversations about these images. To harness the potential of VLMs in the RS domain, existing studies have primarily focused on two approaches for data curation: rule-based methods and synthetic data generation using established VLMs. For rule-based approaches, RS5M (zhang2023rs5m, ) explores the use of binary RS image classifiers to select RS image-text pairs from open-source general vision-language datasets. Similarly, RemoteCLIP (liu2024remoteclip, ), Skyscript (wang2024skyscript, ), and SkysenseGPT (luo2024skysensegpt, ) implement manually designed text templates or relation graphs for data construction. In parallel, studies such as VHM (pang2024vhmversatilehonestvision, ), LHRS-Bot (muhtar2024lhrs, ), LHRS-Bot-Nova (li2024lhrs, ), and GeoChat (kuckreja2024geochat, ) leverage flagship VLMs like GPT4 or Gemini for synthetic vision-language data curation. While these works have advanced the development of RS-specific VLMs, our practical applications reveal that data quality remains suboptimal, motivating the development of automated data selection methods.
2.2 RS VLMs
The remarkable success of VLMs in understanding images and engaging in complex tasks (lu2024omniparser, ; pan2024vlp, ) has inspired the development of specialized VLMs for RS image interpretation. Several approaches have focused on adapting CLIP (radford2021learning, ) for the RS domain. RemoteCLIP (liu2024remoteclip, ), Skyscript (wang2024skyscript, ), and RS5M (zhang2023rs5m, ) fine-tuned CLIP using carefully curated RS image-caption pairs, enabling robust RS zero-shot classification and image-text retrieval capabilities. For larger VLMs, GeoChat (kuckreja2024geochat, ) pioneered the finetuning of LLaVA (liu2024visual, ) with RS-specific vision instruction datasets. Subsequent works have expanded this paradigm through various strategies: incorporating volunteer geographic information (VGI) (muhtar2024lhrs, ), integrating high-quality vision-language data (li2024lhrs, ; pang2024vhmversatilehonestvision, ), modeling relational graphs (luo2024skysensegpt, ), implementing ensembled vision encoders for vision-centric designs (zhang2024earthgpt, ), improving temporal understanding of RS images (irvin2024teochat, ), and scaling data volume with significantly smaller encoder-decoder architectures (yao2025falconremotesensingvisionlanguage, ). The versatile and successful applications of RS VLMs have directed our attention to the fundamental building blocks underlying these models: data quality and utilization. Rather than proposing new architectural designs, our work focuses on addressing how to improve data quality and enhance VLM performance within existing frameworks.
2.3 Parameter-Wise Data Selection
As data volume continues to expand and data complexity evolves, quality control has become increasingly challenging. Beyond traditional rule-based data selection methods (soldaini2024dolma, ; wenzek2019ccnet, ), researchers have shifted focus toward parameter-wise selection approaches for more sophisticated data curation. In the language domain, methods such as QaRater (wettig2024qurating, ) and LESS (xia2024less, ) explore training dedicated scoring models or leveraging model gradients to select high-quality data for LLM pretraining and instruction tuning. Similarly, for vision-language data, Zhang et al. (zhang2023filter, ) investigated scoring models for curating image-text pairs. The model selection approach has been widely adopted in recent state-of-the-art VLMs, including Llama-3.2-Vision (grattafiori2024llama, ), DeepseekVL2 (wu2024deepseek, ), and Qwen2VL (wang2024qwen2, ), all of which implement scoring models for data filtering. Despite these advances, there is a notable absence of open-source scoring models specifically designed to evaluate the quality of RS vision-language data. In this work, we develop and release the first open-source scoring model tailored for RS vision-language data quality assessment.
@@ 3 Method
Figure 2: Pipeline for generating pairwise preference datasets and the training/application of our ScoreRS model. $I_{i}\in\mathcal{I}$ represents a RS image, and $T_{i}\in\mathcal{T}$ represents an image caption, question, or conversation associated with the image
@@
Formally, we define a scoring function $f_{\theta}:\mathcal{T}\times\mathcal{I}\rightarrow\mathbb{R}$, parameterized by $\theta$, where $\mathcal{T}$ and $\mathcal{I}$ denote the text and image spaces, respectively. Given an image-text pair $(I,T)\in\mathcal{I}\times\mathcal{T}$, which may consist of either a simple caption or a multi-turn conversation associated with the image, the model outputs a scalar quality score $s=f_{\theta}(I,T)$. The scoring function is trained to assign higher scores to better-aligned image-text pairs $(I,T)$ through pairwise preference learning. For each image $I$, we collect pairs of text $(T^{+},T^{-})$ where $T^{+}$ is preferred over $T^{-}$. This preference dataset is formally defined as $\mathcal{D}={(I_{i},T^{+}_{i},T^{-}_{i})}_{i=1}^{|\mathcal{D}|}$. The pairwise preferences are modeled using the Bradley-Terry model (bradley1952rank, ), which defines the probability of $T^{+}$ being preferred over $T^{-}$ given image $I$ as:
$p(T^{+}\succ T^{-}|I)=\frac{\exp(f_{\theta}(I,T^{+}))}{\exp(f_{\theta}(I,T^{+}))+\exp(f_{\theta}(I,T^{-}))}.$ (1)
The parameters $\theta$ are optimized by minimizing the empirical negative log-likelihood loss (ouyang2022training, ; wang2024rovrm, ):
@@
Given the absence of RS-specific vision-language preference datasets, we begin by establishing a framework that defines five key dimensions of textual quality for RS images. This framework characterizes what constitutes high-quality descriptions and question-answer pairs, providing the foundation for our data collection pipeline to construct both image-caption and vision instruction preference pairs (Figure 2). Subsequently, we detail the construction of our scoring function $f_{\theta}$ by leveraging pre-trained VLMs and employing a progressive training strategy.
@@ 3.1 Preference Data Construction
3.2 Quality Dimensions for RS Vision-Language Data
To develop an effective scoring function $f_{\theta}$, we must first establish clear criteria for what constitutes high-quality RS vision-language data. These criteria provide the foundation for constructing our preference dataset and subsequently training our score function.
To this end, we define five critical dimensions that characterize high-quality RS vision-language data: relevance (pang2024vhmversatilehonestvision, ; luo2024skysensegpt, ; xue2024enhanced, ), specificity & detail (li2024lhrs, ; li2024if, ; zhou2025benchmark, ), completeness (chan2023clair, ; zhou2025benchmark, ), clarity & fluency (wang2024skyscript, ), and semantic richness (wen2025rs, ; silva2402large, ). For each dimension, we establish a 5-point scoring system (1=Poor to 5=Excellent) with distinct criteria for both image-caption pairs and instruction samples. Detailed descriptions of these quality dimensions, along with illustrative examples and the complete scoring rubrics, are provided in the Appendix A.
@@ 3.2.1 Image-Caption Preference Dataset
Following established practices in VLM training (wang2024qwen2, ; muhtar2024lhrs, ; pang2024vhmversatilehonestvision, ; liu2024visual, ), we begin by constructing an image-caption preference dataset to train ScoreRS to evaluate caption quality for RS images. Given the geographical variations in RS images (muhtar2024lhrs, ; wang2024skyscript, ), we utilize the LHRS-Align dataset (muhtar2024lhrs, ) as our image source. This dataset contains 1.15M orthorectified RS images from major global urban areas, enabling ScoreRS to learn from diverse geographical contexts worldwide. Prior to generating preferences, we implement a rigorous image deduplication process to ensure image quality and representativeness, resulting in 76K distinctive RS images. The detailed method for this deduplication process is provided in Appendix B.1.
For each deduplicated image, we generated captions using three VLMs: RS-specific LHRS-Bot-Nova (li2024lhrs, ) and general-purpose Qwen2VL-7B (wang2024qwen2, ) and InternVL-2.5-8B (chen2024expanding, ). The rationale for selecting these generation models is detailed in Appendix B.2. We generate captions for each image using same prompts and sampling parameters across all models. The resulting captions are then evaluated by GPT-4o based on our predefined scoring system. For each image, we compute the mean score across five dimensions per caption, designating the highest-scoring as positive ($T^{+}$) and others as negative ($T^{-}$). After removing results with parsing errors, we construct a dataset of 72K image-caption preference pairs.
To validate GPT-4o’s reliability as a preference judge, we conduct a human evaluation study on 1,000 random pairs. Human experts independently annotate these samples to identify positive and negative captions. The results show 92.6% agreement (926 out of 1,000 samples) between human judgments and GPT-4o’s assessments. This high level of agreement validates our choice of GPT-4o as a reliable judge for generating pairwise preferences.
@@ 3.2.2 Vision Instruction Preference Dataset
Vision instruction data consists of conversations about images (liu2024visual, ). To construct a RS-specific vision instruction preference dataset that enables ScoreRS to evaluate responses to diverse user queries, we prioritize collecting a broad spectrum of question types and conversation scenarios. We aggregate vision instruction data from multiple sources: GeoChat (kuckreja2024geochat, ) (306K samples), LHRS-Instruct (muhtar2024lhrs, ) (39.8K samples), and a subset of SkysenseGPT (luo2024skysensegpt, ) (381K samples) as the source of our vision instruction preference dataset. To address potential redundancy across these datasets, we implement a two-stage similarity-based filtering process, resulting in 112K diverse and non-redundant instruction samples. The detailed method for this deduplication procedure is provided in the Appendix B.1.
After deduplication, we categorize conversations into close-ended questions (with definitive answers) and open-ended questions (without verifiable answers). For close-ended questions, we prompt Qwen2VL-7B to generate an answer $\hat{y}$. When $\hat{y}$ differed from the provided answer $y$, we extract the sample, treating $\hat{y}$ as negative target $T^{-}$ and $y$ as positive target $T^{+}$.
@@
For open-ended questions, we use LHRS-Bot-Nova and Qwen2VL-7B to generate answers $\hat{y}_{1}$ and $\hat{y}_{2}$, then prompt Qwen2VL-72B to evaluate these along with the source dataset answer $y$ using predefined dimensions. We designate the higher scoring answer as the positive target ($T^{+}$) and others as negative targets ($T^{-}$). We do not directly consider the answer from the source dataset as $T^{+}$ because many of them are generated by older VLMs and are often too short, meaningless, or contain incorrect statements. To manage costs associated with the large volume of data, we employ Qwen2VL-72B rather than GPT-4o as our evaluator. After filtering out parsing errors and applying human-defined quality rules, we curate a dataset of 26K vision instruction preference pairs. The detailed prompts used in this process are provided in Appendix B.3.2.
@@
While collecting question set $Q$, we notice most sources lack RS application-specific questions (e.g., agricultural and disaster analysis). To enhance ScoreRS’s domain-specific performance, we create a specialized set of five manually crafted questions for each of 12 expert-defined RS image analysis categories. We then randomly sample 35K RS images from our deduplicated 112K image dataset. For each image, we randomly select a question from the manually designed question set and prompt Qwen2.5-13B to rephrase the question to increasing the diversity. LHRS-Bot-Nova and Qwen2VL-7B then generate answers based on these image-question pairs. Finally, GPT-4o evaluates the generated answers, selecting the higher-scoring one as the positive target $T^{+}$ and treating others as negative targets $T^{-}$. This process yield 33K RS-specific vision instruction preference data points after removing entries with parsing errors. The manually designed questions and categories can be found in Appendix B.3.2.
3.3 Training and Application
Training
Our ScoreRS model is initialized with Qwen2VL-7B, with the language head replaced by a linear layer to output a scalar score, following the standard value-head-based reward model (wang2024rovrm, ; dong2024rlhf, ). We do not consider generative score modeling due to efficiency concerns. Similar to the standard recipe for training large VLMs (wang2024qwen2, ; liu2024visual, ), we implement a multi-stage training procedure to gradually train ScoreRS to distinguish better vision-language pair. In the first stage, we train the newly introduced value head using a pure text preference dataset, UltraFeedback (cui2024ultrafeedback, ), to provide a good initialization for the value head. Then, we unfreeze the ViT and value head and train ScoreRS on our image-caption preference dataset to enable ScoreRS to better understand RS images. Finally, we unfreeze the LLM and conduct full-parameter training with our vision instruction preference dataset and the additional RLHF-V (yu2024rlhf, ) dataset. This training approach enables ScoreRS to effectively identify high-quality outputs and assign higher scores to better responses across diverse RS scenarios.
Application
Beyond data selection, we also explore ScoreRS’s usage for RL training and BoN selection. We primarily discuss implementing ScoreRS for RL, as its applications to data selection and BoN selection are straightforward. Our RL framework is based on GRPO (shao2024deepseekmath, ), chosen for its computational efficiency and ease of hyperparameter tuning. RL training methods typically require a reward model to evaluate each action trajectory and update the model parameters to favor outputs with higher scores (i.e., responses more aligned with human preferences). While DeepSeek-R1 (guo2025deepseek, ) demonstrated that rule-based rewards for close-ended questions with verifiable answers are effective for RL training, this approach is insufficient for RS applications. In the RS image understanding domain, most questions are open-ended, such as "Describe the urban development patterns visible in this satellite imagery", requiring nuanced interpretations rather than definitive answers. These questions also usually do not have verifiable answers, as different interpretations can lead to multiple acceptable responses. To address this challenge, we introduce a novel reward method using ScoreRS for evaluating open-ended responses while incorporating rule-based rewards for close-ended questions, creating an approach that handles both question types. Specially, for close-ended questions, we employ a binary reward (0 or 1) based on exact match or intersection over union (IoU) with the ground truth. For open-ended questions, where we have a reference answer $y$ (typically sampled from standard vision instruction datasets), we compute the reward $r$ for a predicted answer $\hat{y}$ using the following formulation:
where $\beta>0$ is a hyperparameter controlling the reward sharpness. We use this reference-based approach because it accelerates learning. The reason behind this approach is that the reference answer serves as a baseline, allowing the policy model to improve upon it by generating responses that receive higher scores. This method focuses on continuous improvement rather than selecting the least problematic option from multiple suboptimal alternatives that may be worse than the reference answer itself. We give our detailed unsuccessful attempts and our reasoning in AppendixC.1F.2.
4 Experiment
We evaluate ScoreRS’s effectiveness across three key applications: (1) vision-language data selection for training VLMs, (2) RL training, and (3) BoN selector. We also analyze the impact of score model’s size, initialization, and training strategies, along with our curated preference dataset quality.
4.1 Vision-Language Data Selection
4.1.1 CLIP Finetuning
Experimental Setting
To validate ScoreRS’s effectiveness for data selection, we score and rank training samples from RemoteCLIP (liu2024remoteclip, ) to finetune CLIP-ViT-L/14 (radford2021learning, ). We compare against three baselines: (1) CLIP without finetuning, (2) CLIP finetuned on the complete dataset, and (3) CLIP finetuned on CLIP-score filtered data. We provide detailed training hyperparameters in Appendix D.3 and a detailed description of evaluation methodology, including datasets and metrics, in Appendix D.3.1. Moreover, we also compare using ScoreRS for data filtering with filtering with more advanced VLMs like SigLIP-2 (tschannen2025siglip, ) and domain specific VLMs like CLIP-LION-RS (wang2024skyscript, ) in Appendix C.3.3.
Main Result
We evaluate the finetuned models on classification and retrieval tasks. Results in Table 1 and Table 2 challenge the "more data yields better performance" assumption for RS-specific VLMs. CLIP-score filtering improves performance by 5% (classification) and 1% (retrieval), while our ScoreRS filtering further advances these gains to 7% and 2% respectively, outperforming both the complete dataset and CLIP-score filtering. These findings confirm our hypothesis that current RS vision-language data are suboptimal and require quality control.
Table 1: Comparison of finetuned CLIP models on classification tasks. Top-1 (@1) and top-5 (@5) classification accuracies are reported.